Calculate the pH of an acetic acid buffer with [A-] = [HA] and the pH of an acetic acid buffer with [A-] = 2[HA]

Determine how to create 20mL of each buffer given the follwing chemicals: 0.1 M acetic acid, 0.1 M NaOH,

,NaCH3CO2, and deionized water.
Answer #1

pH of acidic buffer = pka + log(ch3coona/ch3cooh)

pka of ch3cooh = 4.76

if [A-] = [HA]

   [CH3COONa] = [CH3COOH], both are equal concentrations

let us take, [CH3COONa] = [CH3COOH] = 1 M

pH = 4.76+log(1/1)

pH = 4.76

if [A-] = 2[HA]

   [CH3COONa] = 2*[CH3COOH], both are equal concentrations

let us take, [CH3COONa] = 2 M

             [CH3COOH] = 1 M

pH = 4.76+log(2/1)

pH = 5.061

preparation, if [A-] = [HA]

to create 20mL of buffer with [A-] = [HA]

take both in equal ratio, as the concentration of CH3COOH,CH3COONa are equal.

take 10 ml of 0.1 M CH3COOH and 10 ml of 0.1 M ch3cooNa.


if [A-] = 2[HA]

[CH3COONa] = 2 mol

[CH3COOH] = 1 mol

to create 20mL of buffer with [A-] = 2[HA]

take 1:2 volume ratio of CH3COOH,CH3COONa , as the concentration of CH3COOH,CH3COONa are equal.

SO that,

volume of CH3COOH must take = 20/3 = 6.667 ml

volume of CH3COONa must take = 20-6.67 = 13.33 ml